How many four digits numbers are there with no digit repeated

Grade:11

1 Answers

Arun
6 years ago
Dear Priyanshu
 
1st digit can be filled by 9 ways.(excluding 0)
2nd digit can be filled with 9 fills (including 0)
3rd digit cab be filled by 8 ways
4th digit can be filled by 7 ways.
 
Hence total numbers = 9*9*8*7 = 81* 56 = 4536
 
Regards
Arun (askIITians forum expert)